OpenStudy (anonymous):

please help! > The total distance was 540 miles. Part of the journey was on motorcycle at 40mph, and part was in a car at 60mph. What distance was covered by motorcycle if the total time of the journey was 11hours?

OpenStudy (mathstudent55):

distance = speed * time d = vt Let t = time traveled by motorcycle, then 11 - t = time traveled by car For the motorcycle: dm = vt = 40t For car: dc = vt = 60(11 - t) Total distance = dt + dc 540 = 40t + 60(11 - t) 540 = 40t + 660 - 60t -120 = -20t t = 6 The motorcucle traveled 6 hours at 40mph, so d = vt = 40 * 6 = 240 The distance traveled in the motorcylce was 240 miles